# GUARD: Constructing Realistic Two-Player Matrix and Security Games for Benchmarking Game-Theoretic Algorithms

### v1

@@ 1 Introduction
To fill this gap in benchmarking ability, we introduce GUARD, a flexible framework that constructs realistic matrix and security game instances from open-access data (open-source and publicly-available at https://github.com/CoffeeAndConvexity/GUARD). Drawing on readily available sources such as animal movement records for anti-poaching scenarios and demographic or infrastructure data for critical asset protection, our framework supports both custom game specification and access to a library of ready-made instances, enabling meaningful and reproducible evaluation of game-theoretic algorithms on security-inspired scenarios. The framework provides a more realistic real-world resource allocation and security grounded alternative to existing game generators such as Gamut nudelman2004run , which includes more traditional and classical economic games, or OpenSpiel LanctotEtAl2019OpenSpiel , which offers a wide range of recreational and synthetic benchmarks. Our game instances can be exported in formats compatible with tools like OpenSpiel and Gambit savani2025gambit , and can also provide target values for integration into pursuit-evasion frameworks such as GraphChase zhuang2025solving .
@@ B.1 Schedule-Finding Algorithms
Algorithm 1 and the subroutine given by Algorithm 2 are used to return valid general (multiple target coverages per schedule allowed) schedules per defender resource.
@@
1:Graph $G=(V,E)$, start node $h$, number of timesteps $T$, interdiction threshold $\tau$, dwell cost per timestep $c$
@@
3:$S_{\text{simple}}\leftarrow$ GetSimpleSchedules($G,h,T,\tau$)
@@
10:         $(p,\text{cost},\text{steps})\leftarrow$ GetFullPath($G,h,S,\tau$)
@@
11:              Append $v$ to $p$ $(\tau-1)$ times $\triangleright$ Dwell at $v$
12:              $c\leftarrow c+\tau-1$
@@
The simple schedule case is handled via a closed-form condition: for any target $v$ and home base $h$, if the shortest path $p$ from $h$ to $v$ satisfies $t\geq 2|p|+d-1$, then target $v$ is considered validly defendable as a singleton schedule.
@@ B.2.1 Best Response Oracles in Normal-Form Double Oracle
As part of the double oracle algorithm for solving zero-sum security games, the normal-form defender best response subroutine DefenderBRNF solves a mixed-integer program to determine the optimal joint strategy for the defender team, given a fixed distribution over attacker strategies. This best response accounts for spatial movement constraints, home base assignments, and an interdiction threshold $\delta$ specifying the number of time steps a defender must occupy a target to successfully interdict it.
Let $G=(V,E)$ be a directed graph over which defender movement is defined, and let $\mathcal{T}$ denote the number of timesteps in the game. The defender team consists of $D$ mobile units, each with a valid home base set $H_{d}\subseteq V$. Let $T\subseteq V$ be the full set of targets, and let $T_{a}\subseteq T$ denote the subset of targets currently selected by the attacker. Each $t\in T_{a}$ is associated with a target value $V_{t}$ and a probability of attack $P_{t}$.
@@
For our Normal-Form Double Oracle algorithm, we define the AttackerBR as selecting the $k$ targets with highest expected attacker value, where utility is discounted by the probability $q_{t}$ that target $t$ is interdicted under the current defender strategy $x_{d}$. Formally, the attacker solves:
$\textsc{AttackerBR}(x_{d})=\underset{S\subseteq T,\,|S|=k}{\arg\max}\sum_{t\in S}V^{a}_{t}\cdot(1-q_{t})$
where $T$ is the set of all targets, $V^{a}_{t}$ is the attacker value for target $t$, and $q_{t}$ is the interdiction probability of $t$ induced by defender strategy $x_{d}$. The attacker selects the top-$k$ targets with the highest discounted expected utility.
@@ B.2.2 Best Response Oracles in Schedule-Form Double Oracle
In schedule-form security games, each defender resource selects a schedule—a subset of targets it can reach and cover—subject to mobility and dwell-time constraints. Given a fixed mixed strategy of the attacker, the defender’s best response selects one schedule per resource to maximize expected utility. Let $\mathcal{R}$ denote the set of defender resources, $T$ the set of targets, and $\mathcal{S}_{r}\subseteq 2^{T}$ the set of feasible schedules for defender resource $r\in\mathcal{R}$. Let $w_{t}$ denote the expected utility of defending target $t$, computed from the current attacker strategy.
We define binary variables $x_{r,i}\in\{0,1\}$ indicating selection of the $i^{\text{th}}$ schedule in $\mathcal{S}_{r}$, and $g_{t}\in\{0,1\}$ indicating whether target $t$ is covered. The defender best response solves:
@@
$\textsc{AttackerBR}(x_{d})=\arg\max_{t\in T}\sum_{j}D_{d}(j)\cdot\begin{cases}V^{a}_{\text{c},t}&\text{if }t\in S_{j}\\ V^{a}_{\text{uc},t}&\text{otherwise}\end{cases}$
@@ B.3 Time Constrained Depth First Search
To enumerate the full set of valid movement actions for a team of mobile units on a graph, we employ a time-constrained depth-first search (DFS). This recursive procedure explores all paths of a fixed length $\mathcal{T}$, starting from allowed initial nodes and optionally constrained by required end nodes or a return-to-start condition. Waiting behavior is supported by including self-loops when allowed. For multiple units, individual path sets are generated and combined via Cartesian product, then reformatted into timestep-major action matrices for use in normal-form strategy construction.
@@ Appendix C Framework Details
This Appendix details the the the GUARD game class hierarchy in greater detail than the main text. For exact implementation details, please visit our open-source Git repository for this project at https://github.com/CoffeeAndConvexity/GUARD.
@@ C.1 Graph Game Parameters
  • Unit counts for each player: number of moving/stationary attackers and defenders.
@@
  • A set of Target objects, each with attacker and defender values.
  • Additional game constraints, such as waiting permissions (allow_wait) and forced return to origin (force_return).
@@
The generate_actions(player) method constructs a valid joint action space for either the attacker or defender, handling both moving and stationary units. Moving unit actions are computed using a constrained depth-first search that respects home bases, end nodes, waiting rules, and return constraints. Stationary attacker actions select from subsets of the target nodes (or include None), while stationary defender actions are generated via Cartesian products of valid placements. Combined action paths are returned in timestep-major form as $\mathcal{T}\times N$ arrays.
@@
The evaluate_actions(defender_action, attacker_action) method simulates interactions between defender and attacker paths using the game’s interdiction protocol. Attacker utilities are calculated by summing the values of targets reached without being interdicted. Defender values are represented as the negation of attacker scores in zero-sum settings or separately in general-sum settings. Interdiction for moving attackers is radius-based and computed at each timestep; interdiction for stationary attackers is governed by a defense time threshold requiring $\delta$ visits.
@@
The evaluate_actions_general method extends utility computation by incorporating defender costs proportional to the number of steps taken. This enables richer modeling of path efficiency and asymmetric incentives.
Extensibility.
The modular structure of the Graph Game class enables its reuse across domain-specific instances (e.g., Green Security Games, Infrastructure Security Games).
@@ C.2 Security Game Parameters
  • An InterdictionProtocol object, which governs coverage rules based on dwell time thresholds and graph-based distances.
@@
In schedule-form mode, each defender is assigned a set of feasible schedules (i.e., subsets of targets reachable within the time horizon $\mathcal{T}$ while satisfying interdiction dwell time requirements and optional return constraints). These schedules are generated via a recursive backtracking procedure with pruning (see Algorithm 1) and then deduplicated.
@@
This abstraction isolates core SSG functionality while remaining modular enough to support domain-specific modeling in higher layers, including target value scaling, defender mobility constraints, and payoff asymmetries.
@@ C.3 Domain-Specific Game Parameters
The GreenSecurityGame and InfraSecurityGame classes extend the SecurityGame layer to support context-specific initialization, input data processing, and target construction. Both game types ultimately instantiate and return a configured SecurityGame object, but differ in how graphs, targets, and values are derived.
@@
  • interdiction_protocol: Interdiction rules (e.g., dwell time threshold).
@@
GSGs define their graph using a custom spatial grid over a bounding box. Each cell contains animal tracking data and is scored using either density (raw frequency of observations) or centroid (K-means clustering with cluster weights). Optional escape-line proximity adjustments are applied to model attacker incentives for exit routes. Targets are created using the cell center, population density, and animal value scaling factors.
@@
ISGs use OpenStreetMap street networks and power infrastructure feature data to construct real-world graphs. Nodes are street intersections, and targets are mapped from power infrastructure features to the closest node. Target values are computed using population estimates (from Census block data) and infrastructure type multipliers.
@@
  • power_df: DataFrame of infrastructure features with coordinates and types.
@@
  • infra_weights: Dictionary of weights per infrastructure type (e.g., substation, power_tower).
@@
  • attacker_feature_value, defender_feature_value: Scaling factors for general-sum target utilities.
@@
If general-sum mode is enabled, the system computes asymmetric target payoffs:
  • attacker_penalty_factor: Scales down attacker value when target is interdicted.
  • defender_penalty_factor: Scales down defender value when interdiction is successful.
@@ C.7 Potential Negative Societal Impacts
To mitigate such risks, we introduce several precautions. First, researchers and practitioners using the framework should refrain from publicly releasing sensitive configuration files, particularly those derived from real defense operations or infrastructure schematics (all data we utilize is publicly available and does not compromise any specific assets). Second, future development can incorporate access controls for datasets and models flagged as high-risk. Finally, if usage/develpoment is widespread, collaboration with domain experts can ensure responsible deployment and use of the GUARD framework, particularly in settings involving law enforcement, conservation, or urban infrastructure.
@@ C.8 Pre-Defined Game Instance Details
Table 2: Pre-generated GSG instances for Lobéké National Park. $\mathcal{T}$ = Timesteps, A = Attackers, D = Defenders, $\delta$ = Defense Time Threshold, FR = Force Return, S = Nash Support, Size = Utility matrix dimensions ($|A_{d}|\times|A_{a}|$), T = True, F = False.
@@ D.4 Stackelberg Experiment Parameters
  • Defender step cost: 1.17 (realistic labor + fuel cost per km)
@@
All ISG Stackelberg experiments used the same Chinatown-area power grid graph with a fixed escape point at Brooklyn Bridge ((40.7124, -74.0049)). Defenders started at the same 5 precinct locations used in all ISG experiments.
@@
Defender elephant value of \$22,966 reflects the one year eco-tourism value of an elephant. The attacker value is computed as an elephant’s 2 tusks $\times$ 100 pound tusk average $\times$ \$26/kg on the Cameroonian black-market is roughly \$2,350  [41, 17]. These values serve to normalize the payoffs. Because they are constant scalers on target values, they do not tangibly impact the equilibria. However, the other general sum factors like path costs and escape proximities do impact the equilibria.
@@ D.5 Schedule Randomization
Schedules were randomized by assigning each defender the same number of schedules as in the real instance, with each schedule containing randomly selected targets equal to the ceil of the average real schedule length.
D.6 Infra Weights (for ISGs)
Listing 5: Infrastructure Security Game feature base weights, manually assigned based on qualitative assessments of each feature’s relative criticality. While the default values serve as reasonable approximations, the framework is designed to allow users to adjust these weights to reflect application-specific priorities or expert-provided valuations.

### v2

@@ 1 Introduction
To fill this gap in benchmarking ability, we introduce GUARD, [^1] Code open-sourced and publicly-available at https://github.com/CoffeeAndConvexity/GUARD. a flexible framework that constructs realistic matrix and security game instances from open-access data. Drawing on readily available sources such as animal movement records for anti-poaching scenarios and demographic or infrastructure data for critical asset protection, our framework supports both custom game specification and access to a library of ready-made instances, enabling meaningful and reproducible evaluation of game-theoretic algorithms on security-inspired scenarios. The framework provides a more realistic real-world resource allocation and security grounded alternative to existing game generators such as Gamut nudelman2004run , which includes more traditional and classical economic games, or OpenSpiel LanctotEtAl2019OpenSpiel , which offers a wide range of recreational and synthetic benchmarks. Our game instances can be exported in formats compatible with tools like OpenSpiel and Gambit savani2025gambit , and can also provide target values for integration into pursuit-evasion frameworks such as GraphChase zhuang2025solving .
@@ B.1 Schedule-Finding Algorithms
Algorithm 1 and the subroutine given by Algorithm 2 are used to return valid general schedules (multiple target coverages per schedule allowed) per defender resource.
@@
1:Graph $G=(V,E)$, start node $h$, number of timesteps $\mathcal{T}$, defense time threshold $\delta$, path cost per edge traveled $c$
@@
3:$S_{\text{simple}}\leftarrow$ GetSimpleSchedules($G,h,\mathcal{T},\delta$)
@@
10:         $(p,\text{cost},\text{steps})\leftarrow$ GetFullPath($G,h,S,\delta$)
@@
11:              Append $v$ to $p$ $(\delta-1)$ times $\triangleright$ Dwell at $v$
12:              $c\leftarrow c+\delta-1$
@@
The simple schedule case is handled via a closed-form condition: for any target $t$ and home base node $h$, if the shortest path $p$ from $h$ to $t$, in a game with defense time threshold $\delta$, satisfies $\mathcal{T}\geq 2|p|+\delta-1$, then target $t$ is considered validly defendable as a singleton schedule, as this condition implies that the game length $\mathcal{T}$ is long enough for a defender resource to travel to it from its home base, wait the required number of steps to meet the defense time threshold, and return to home base.
@@ B.2.1 Best Response Oracles in Normal-Form Double Oracle
As part of the double oracle algorithm for solving zero-sum security games, the normal-form defender best response subroutine DefenderBRNF solves a mixed-integer program to determine the optimal joint strategy for the defender, given a fixed distribution over attacker actions. This best response accounts for spatial movement constraints, home base assignments, and a defense time threshold $\delta$.
Let $G=(V,E)$ be a directed graph over which defender movement is defined, and let $\mathcal{T}$ denote the number of timesteps in the game. The defender team consists of $D$ mobile resources, each with a valid home base set $H_{d}\subseteq V$. Let $T\subseteq V$ be the full set of targets, and let $T_{a}\subseteq T$ denote the subset of targets currently selected by the attacker. Each $t\in T_{a}$ is associated with a target value $V_{t}$ and a probability of attack $P_{t}$.
@@
For our Normal-Form Double Oracle algorithm, we define the AttackerBRNF as selecting the $k$ targets with highest expected attacker value, where utility is discounted by the probability $q_{t}$ that target $t$ is interdicted under the current defender strategy $x_{d}$. Formally, the attacker solves:
$\textsc{AttackerBRNF}(x_{d})=\underset{S\subseteq T,\,|S|=k}{\arg\max}\sum_{t\in S}V_{t}\cdot(1-q_{t})$
where $T$ is the set of all targets, $V_{t}$ is the value for target $t$, and $q_{t}$ is the interdiction probability of $t$ induced by defender strategy $x_{d}$. The attacker selects the top-$k$ targets with the highest discounted expected utility.
@@ B.2.2 Best Response Oracles in Schedule-Form Double Oracle
In schedule-form security games, each defender resource selects a schedule—a subset of targets it can reach and cover—subject to mobility and defense time threshold constraints. Given a fixed mixed strategy of the attacker, the defender’s best response selects one schedule per resource to maximize expected utility. Let $\mathcal{R}$ denote the set of defender resources, $T$ the set of targets, and $\mathcal{S}_{r}\subseteq 2^{T}$ the set of feasible schedules for defender resource $r\in\mathcal{R}$. Let $w_{t}$ denote the expected utility of defending target $t$, computed from the current attacker strategy.
We define binary variables $x_{r,i}\in\{0,1\}$ indicating selection of the $i^{\text{th}}$ schedule in $\mathcal{S}_{r}$, and $g_{t}\in\{0,1\}$ indicating whether target $t$ is covered. The defender best response DefenderBRSF solves:
@@
$\textsc{AttackerBRSF}(x_{d})=\arg\max_{t\in T}\sum_{j}D_{d}(j)\cdot\begin{cases}V^{a}_{\text{c},t}&\text{if }t\in S_{j}\\ V^{a}_{\text{uc},t}&\text{otherwise}\end{cases}$
@@ B.3 Time Constrained Depth First Search
To enumerate the full set of valid movement actions for a team of mobile units on a graph, we employ a time-constrained depth-first search (DFS). This recursive procedure explores all paths of a fixed length $\mathcal{T}$, starting from allowed initial nodes and optionally constrained by required end nodes or a return-to-start condition. Waiting behavior is supported by including self-loops when allowed. For multiple units, individual path sets are generated and combined via Cartesian product, then reformatted into time-expanded action matrices for use in normal-form action space construction.
@@ Appendix C Framework Details
This Appendix details the the the GUARD game class hierarchy in greater detail than the main text. For exact implementation details, please visit our public Git repository for this project at https://github.com/CoffeeAndConvexity/GUARD.
@@ C.1 Graph Game Parameters
  • Resource counts for each player: number of moving/stationary attackers and defenders.
@@
  • A set of Target objects, each with attacker and defender value attributes.
  • Additional game constraints: waiting permissions (allow_wait) and forced return to origin for moving resources (force_return).
@@
The generate_actions(player) method constructs a valid joint action space for either the attacker or defender, handling both moving and stationary resources. Moving resource actions are computed using a constrained depth-first search that respects home bases, end nodes, waiting rules, and return constraints (see Appendix B.3). Stationary attacker actions select from subsets of the target nodes (or include None), while stationary defender actions are generated via Cartesian products of valid placements. Combined action paths are returned in timestep-major form as $\mathcal{T}\times N$ arrays.
@@
The evaluate_actions(defender_action, attacker_action) method simulates interactions between defender and attacker paths using the game’s interdiction protocol. Attacker utilities are calculated by summing the values of targets reached without being interdicted. Defender values are represented as the negation of attacker scores in zero-sum settings or separately in general-sum settings. Interdiction for moving attackers is radius-based and computed at each timestep; interdiction for stationary attackers is governed by a defense time threshold requiring $\delta$ timesteps spent by defender resources at attacker-selected targets.
@@
The evaluate_actions_general method extends utility computation by incorporating defender path costs. This enables richer modeling of path efficiency and asymmetric incentives.
@@ C.2 Security Game Parameters
  • An InterdictionProtocol object, which governs coverage rules based on defense time thresholds and graph-based distances.
@@
In schedule-form mode, each defender is assigned a set of feasible schedules (i.e., subsets of targets reachable within the time horizon $\mathcal{T}$ while satisfying interdiction dwell time requirements and optional return constraints). These schedules are generated via a recursive backtracking procedure (see Algorithm 1) and then deduplicated.
@@
This abstraction isolates core Stackelberg security-game functionality while remaining modular enough to support domain-specific modeling in higher layers, including target value scaling, defender mobility constraints, and payoff asymmetries.
@@ C.3 Domain-Specific Game Parameters
The GreenSecurityGame and InfraSecurityGame classes extend the SecurityGame layer to support input data processing and target construction. Both game types ultimately instantiate and return a configured SecurityGame object, but differ in how graphs and targets are derived.
@@
  • interdiction_protocol: Interdiction rules (e.g., defense time threshold).
@@
As detailed in the main text, GSGs define their graph using a custom spatial grid over a bounding box. Each cell contains animal tracking data and is scored using either density (raw frequency of observations) or centroid (K-means clustering with cluster weights). Optional escape-line proximity adjustments are applied to model attacker incentives for exit routes. Targets are created using the cell center, population density, and animal value scaling factors.
@@
As mentioned in the main text, ISGs use OpenStreetMap street networks and power infrastructure feature data to construct real-world graphs. Nodes are street intersections, and targets are mapped from power infrastructure features to the closest node. Target values are computed using population estimates (from Census block data) and infrastructure type multipliers.
@@
  • infra_df: DataFrame of infrastructure features with coordinates and types.
@@
  • infra_weights: Dictionary of weights per infrastructure type (e.g., medical clinic, power_tower).
@@
  • attacker_feature_value, defender_feature_value: Scaling weights for general-sum target utilities, serves to normalize payoffs.
@@
If general-sum mode is enabled, the system allows asymmetric target coverage scaling factors:
  • attacker_penalty_factor: Scales down (positive) gained attacker value when a target is covered by a defender schedule.
  • defender_penalty_factor: Scales down (negative) incurred defender value when a target is covered by a defender schedule.
@@ C.7 Potential Negative Societal Impacts
To mitigate such risks, we introduce several precautions. First, researchers and practitioners using the framework should refrain from publicly releasing sensitive configuration files, particularly those derived from real defense operations or infrastructure schematics (all data we utilize is publicly available and does not compromise any specific assets). Second, future development can incorporate access controls for datasets and models flagged as high-risk. Finally, if usage/development becomes widespread, collaboration with domain experts can ensure responsible deployment and use of the GUARD framework, particularly in settings involving law enforcement, conservation, or urban infrastructure.
@@ C.8 Pre-Defined Game Instance Details
Table 2: Pre-generated GSG instances for Lobéké National Park. C = Num. Elephant Cluster Targets, Grid = Game gridworld dimensions $\mathcal{T}$ = Timesteps, A = Attackers, D = Defenders, $\delta$ = Defense Time Threshold, FR = Force Return, S = Nash Support, Size = Utility matrix dimensions ($|A_{d}|\times|A_{a}|$), T = True, F = False.
@@ D.4 Stackelberg Experiment Parameters
  • Defender step cost: 1.17 (realistic labor, fuel, and equipment cost per km)
@@
All ISG Stackelberg experiments used the same Chinatown-area power grid graph with a fixed escape point at Brooklyn Bridge (40.7124, -74.0049). Defenders were allowed to start at any of the same 5 precinct locations used in all ISG experiments.
@@
Defender elephant value of \$22,966 reflects the one year eco-tourism value of an elephant. The attacker value is computed as an elephant’s 2 tusks $\times$ 100 pound tusk average $\times$ \$26/kg on the Cameroonian black-market is roughly \$2,350  [41, 17]. These values serve to normalize the payoffs. Because they are constant scalers on target values, they do not tangibly impact the equilibria. However, the other general sum factors like path costs and escape proximities do have material impact on the equilibria as they decouple target values nonlinearly.
@@ D.5 Schedule Randomization
Schedules were randomized by assigning each defender the same number of schedules as in the real instance, with each schedule containing as many randomly selected targets equal to the ceil of the average real schedule length.
D.6 Additional Experiment Results
Below are additional iterative and Stackelberg experiment results from our realistic security game instances, illustrating the strategic complexity and large relative support sizes characteristic of settings with enhanced realism through the GUARD framework. All home base, scaling factors, and general sum (for Stackelberg) parameter settings are equal to those used in the experiments in the main text and detailed in Appendices D.3 and D.4.
Form C Grid $\mathcal{T}$ A D $\delta$ FR S Time (s) DO Size Iters
NFG 12 $10\times 10$ 10 1 3 1 F 12 98.25 $45\times 12$ 46
NFG 12 $9\times 9$ 10 3 3 1 F 12 40.79 $37\times 41$ 41
NFG 12 $8\times 8$ 10 2 3 1 T 11 72.71 $32\times 31$ 33
NFG 11 $8\times 8$ 9 2 3 2 T 11 32.93 $27\times 27$ 28
NFG 10 $7\times 7$ 8 2 3 2 F 10 23.07 $28\times 27$ 29
SFG 12 $8\times 8$ 9 1 2 2 T 9 0.11 $31\times 10$ 35
SFG 10 $10\times 10$ 7 1 3 2 T 9 0.17 $23\times 9$ 24
Table 5: Results from GSG iterative double oracle experiments. Columns: C = Clusters, $\mathcal{T}$ = Timesteps, A = Attackers, D = Defenders, $\delta$ = Defense Time Threshold, FR = Force Return, S = Double Oracle Support, DO Size = Final action space dimensions of the DO subgame ($|A_{d}|\times|A_{a}|$), Iters = Iterations to Converge.
Form $\mathcal{T}$ A D $\delta$ FR S Time (s) DO Size Iters
NFG 9 3 3 2 F 16 671.05 $27\times 33$ 33
NFG 10 2 3 2 F 15 1661.28 $31\times 32$ 32
NFG 9 3 3 1 F 15 2303.54 $41\times 43$ 43
NFG 10 1 3 2 F 14 1494.30 $47\times 17$ 47
NFG 8 2 3 1 F 14 162.24 $27\times 28$ 28
SFG 9 3 2 1 T 12 0.28 $37\times 13$ 36
SFG 6 3 1 1 T 10 0.16 $28\times 12$ 28
Table 6: Results from ISG iterative double oracle experiments.
Form C Grid $\mathcal{T}$ A D $\delta$ FR S Time (s) Def. Utility
SFG $\to$ NFG 10 $7\times 7$ 9 1 3 2 T 9 23.62 -0.133
SFG $\to$ NFG 10 $8\times 8$ 8 1 3 1 T 9 5.08 -0.150
SFG $\to$ NFG 10 $10\times 10$ 7 1 2 1 T 9 0.19 -0.230
SFG $\to$ NFG 10 $7\times 7$ 9 1 1 1 T 8 0.03 -0.285
SFG $\to$ NFG 8 $8\times 8$ 9 1 2 2 T 8 0.06 -0.204
Table 7: SSE experiment results for GSGs (general schedules, general sum, multiple LP SSE solver). SFG $\to$ NFG indicates normal form matrix expansion. Columns: C = Clusters, $\mathcal{T}$ = Timesteps, A = Attackers, D = Defenders, $\delta$ = Defense Time Threshold, FR = Force Return, S = Support Size, Def. Utility = Defender max utility at equilibrium.
Form $\mathcal{T}$ A D $\delta$ FR S Time (s) Def. Utility
SFG $\to$ NFG 7 1 3 1 T 14 $8$0.408
SFG $\to$ NFG 9 1 3 2 T 13 762.70 -0.408
SFG $\to$ NFG 8 1 3 1 T 13 2856.51 -0.404
SFG $\to$ NFG 7 1 3 2 T 12 113.75 -0.422
SFG $\to$ NFG 6 1 3 1 T 10 224.72 -0.421
Table 8: SSE experiment results for ISGs (general schedules, general sum, multiple LP SSE solver).
D.7 Infra Weights (for ISGs)
Listing 5: Infrastructure Security Game feature base weights, manually assigned based on qualitative assessments of each feature’s relative criticality. While the default values serve as reasonable approximations, the framework is designed to allow users to adjust the weights they use to reflect application/domain-specific valuations.